OpenStudy (anonymous):

find the distance between the two points G(3,0), H(8,10)

jimthompson5910 (jim_thompson5910):

d = sqrt((x2-x1)^2+(y2-y1)^2) d = sqrt((8-3)^2+(10-0)^2) d = sqrt((5)^2+(10)^2) d = sqrt(25+100) d = sqrt(125) d = 5*sqrt(5) So exact distance between the two points is 5*sqrt(5) units.

OpenStudy (anonymous):

 from 3 to 8 is 5 to the right, from 0 to 10 is 10 up so you get \[d^2=5^2+10^2=125\] and so \[d=\sqrt{125}\]

OpenStudy (anonymous):

same as jimthompson answer with distance formula.